The sum of two integers is -8 and their product is -20. What are the intergers?
Answer by checkley79(3341)

X+Y=-8
XY=-20
X=-8-Y
(-8-Y)Y=-20
-8Y-Y^2=-20
Y^2+8Y-20=0
(Y+10)(Y-2)=0
Y+10=0
Y=-10 ans.
Y-2=0
Y=2 ans.
X-10=-8
X=10-8
X=2 ans.
X+2=-8
X=-8-2
X=-10 ans.
PROOF:
-2*10=-20
-20=-20
10*-2=-20
-20=-20